OpenStudy (anonymous):

Mr. Fields spend 6 hours on a trip out into the country and back. he walked out at the rate of 4 mph and walked back at the rate of 2 mph. how far out into the country did he go?

OpenStudy (inkyvoyd):

2*4+4*2=16

hero (hero):

d = rt d = 4*6 d = 24 mi

OpenStudy (anonymous):

find two expressions for the time put d = distance then time going is \(\frac{d}{4}\) and time returning is \(\frac{d}{2}\) total time is \[\frac{d}{4}+\frac{d}{2}=6\] solve this for d

OpenStudy (anonymous):

\[\frac{6d}{8}=6\] \[6d=6\times 8\] \[d=8\]

OpenStudy (inkyvoyd):

Whoops, I forgot to divide my number by two. Sorry!

OpenStudy (anonymous):

it's ok..

hero (hero):

@satellite, I think the question only wanted how far into the country Mr. Fields went.

OpenStudy (anonymous):

yes, that is what i solved for

OpenStudy (anonymous):

@Hero 6 is the total time there and back, not the time going we can check that the answer is correct. 8 miles there, and \[\frac{8}{4}+\frac{8}{2}=2+4=6\] so six hours total

hero (hero):

I've done problems similar to this. It means 6 hours there, and 6 hours back. If it means something other than that, then these problems are inconsistent.

hero (hero):

Trust me, I looked it up personally and other d = rt problems had that same "6 hours there and back" type of question

hero (hero):

I don't like the wording of them because it can be mis-interpreted.